# Solving High-dimensional Inverse Problems Using Amortized Likelihood-free Inference with Noisy and Incomplete Data

### v1

@@ 3.2 Summary network
The summary network serves as a critical preprocessing step, which converts the $u$ measurements into a fixed, low-dimensional input (summary statistics) to the conditional NF network. This step is important because likelihood-free methods relies on the summary statistics capturing sufficient information about the measurements to ensure the inference quality. While summary statistics have traditionally been manually crafted based on domain knowledge, such choices may be unreliable, potentially losing important information and adversely impacting the inference process.
The adoption of a summary network offers several advantages, particularly in reducing the data dimensionality and improving the efficiency of NF training. First, it alleviates the computational burden by compressing high-dimensional raw data into a more manageable form, which enhances both training speed and precision. Secondly, preprocessing through the summary network helps remove redundancies and noise inherent in the raw data, which ensures that only the most relevant information is passed on to the inference stage. Finally, the summary network automates the extraction of informative statistics directly from the data, eliminating the need for manual feature selection [24].
Mathematically, summary network consists of the parameterized transformation
@@
where $\hat{\mathbf{u}}$ represents the learned summary statistics from the raw data $\mathbf{u}_{o,1:k}$, which will be utilized as input to the conditional NF. The architecture of the summary network is chosen based on the characteristics of the observation data. For instance, bidirectional long short-term memory (LSTM) networks are especially suited for time-series data, enabling the effective handling of sequential measurements that exhibit long memory and non-linear characteristics. Here, we use a 1D convolutional neural network (CNN) adapted to learn summary statistics of temporal responses.
@@
$\mathcal{L}(\bm{\alpha},\bm{\beta})=\frac{1}{M}\sum_{i=1}^{M}\left\{\frac{1}{2}\left[\mathbf{h}_{\bm{\alpha}}\left(\bm{\theta}^{(i)};\mathcal{\bm{F}}_{\bm{\beta}}\left(\mathbf{u}_{o,1:k}^{(i)}\right)\right)\right]^{2}-\log\left|\det\left(\frac{\partial\mathbf{h}_{\bm{\alpha}}\left(\bm{\theta}^{(i)};\mathcal{\bm{F}}_{\bm{\beta}}\left(\mathbf{u}^{(i)}_{o,1:k}\right)\right)}{\partial\bm{\theta}}\right)\right|\right\}$ (23)
@@ 3.3 Inference network
The core element of NFs is invertible transformations. In our study, we employ two families of parameterized invertible transformations: conditional Invertible Neural Networks (cINNs) and conditional Neural Spline Flows (cNSFs) based on conditional affine coupling layers (cACLs) and conditional spline layers (cSLs), respectively. In this section, we present an inference network combining cACLs and cSLs that is capable of bidirectional mapping between posterior and latent distributions. This dual mapping capability is critical for enabling the forward transformation as well as the reverse propagation.
3.3.1 Conditional invertible neural network (cINN)
Dinh et al. [44] introduced affine coupling layers (ACLs) as a technique to create invertible maps with cheap-to-evaluate Jacobian determinants. Each ACL operates by partitioning the input vector $\bm{\theta}$ into two segments, $\bm{\theta}_{1}$ and $\bm{\theta}_{2}$, and similarly dividing the output vector $\mathbf{z}$ into $\mathbf{z}_{1}$ and $\mathbf{z}_{2}$. This division allows us to define invertible transformations in terms of embedded subnetworks, denoted as $\mathbf{s}_{1}(\cdot)$, $\mathbf{s}_{2}(\cdot)$, $\mathbf{t}_{1}(\cdot)$, and $\mathbf{t}_{2}(\cdot)$. These subnetworks, which can be chosen as arbitrary neural network architectures—like fully connected networks—are not required to be invertible. With the observed summary data $\hat{\mathbf{u}}$ serving as an additional conditioning input, the forward transformations, as shown in Fig. 3(a) are formulated as
@@ 3.3.1 Conditional invertible neural network (cINN)
Figure 3: Illustration of transformations in cACL
The forward transformation can be readily inverted. Given $\mathbf{z}=(\mathbf{z}_{1},\mathbf{z}_{2})$, the inversion is accomplished by inversely transforming through the conditional affine coupling layer (cACL). The inverse transformation, as shown in Fig. 3(b), is formulated as follows
@@
$\bm{\theta}_{1}=(\mathbf{z}_{1}-\mathbf{t}_{2}(\mathbf{\bm{\theta}}_{2},\hat{\mathbf{u}}))\odot\exp(-\mathbf{s}_{2}(\mathbf{z}_{2},\hat{\mathbf{u}}))$ (27)
Eqs. 24-27 are essential in guaranteeing the tractability of the Jacobian determinant of the transformation, like in Eq. 14. Within an ACL, the Jacobian matrix is usually triangular, leading to cost-effective evaluation of its determinant.
@@
$\frac{\partial(\bm{\theta}_{1},\bm{\theta}_{2})}{\partial(\mathbf{z}_{1},\mathbf{z}_{2})}=\begin{pmatrix}\frac{\partial\bm{\theta}_{1}}{\partial\mathbf{z}_{1}}&\frac{\partial\bm{\theta}_{1}}{\partial\mathbf{z}_{2}}\\ \frac{\partial\bm{\theta}_{2}}{\partial\mathbf{z}_{1}}&\frac{\partial\bm{\theta}_{2}}{\partial\mathbf{z}_{2}}\end{pmatrix}=\begin{pmatrix}\mathbf{I}&0\\ \frac{\partial\bm{\theta}_{2}}{\partial\mathbf{z}_{1}}&\frac{\partial\bm{\theta}_{2}}{\partial\mathbf{z}_{2}}\end{pmatrix}$ (28)
@@
In this block triangular structure, the upper left block is the identity matrix $\mathbf{I}$, since $\bm{\theta}_{1}$ is transformed independently of $\mathbf{z}_{2}$. The 0 in the top-right block reflects the fact that $\bm{\theta}_{1}$ is independent of $\mathbf{z}_{2}$. The lower left block captures the dependence of $\bm{\theta}_{2}$ on $\mathbf{z}_{1}$, while the lower right block represents the transformation of $\bm{\theta}_{2}$ with respect to $\mathbf{z}_{2}$. The determinant of a block triangular matrix is simply the product of the determinants of the diagonal blocks. Since the upper left block is the identity matrix, its determinant is 1, and the overall determinant of the Jacobian matrix is determined by the diagonal terms.
3.3.2 Conditional neural spline flow (cNSF)
@@
Rational-Quadratic Neural Spline Flows (NSF) are designed to model complex distributions with high efficiency and flexibility [45]. NSFs employ monotonic rational-quadratic splines as the fundamental building block for developing invertible functions, referred to as ’spline layers’. Each bin within these layers is characterized by a monotonically increasing rational-quadratic function. Rational-quadratic functions, defined as the ratio of two quadratic polynomials, are not only differentiable but also analytically invertible when restricted to monotonic segments. The knot points refer to the boundaries between consecutive bins in the spline. At each knot point, the spline must satisfy certain conditions, such as continuity in both the function’s value and its derivative, ensuring smooth transitions between the segments. NSF provides more flexibility than traditional quadratic functions because they allow direct control over the values and derivatives at each knot point. A monotonic rational-quadratic spline is designed to map an interval of $[-B,B]$ onto itself. Outside of this range, the transformation defaults to an identity function, thereby introducing linear ‘tails’ to the curve.
Figure 4: Monotonic rational-quadratic transformation (Spline layer)
@@
Fig. 4 illustrates a monotonic rational-quadratic transformation, also known as a spline layer, which is conditioned on the observed data $\hat{\mathbf{u}}$. Splines are applied component-wise, meaning that each dimension is transformed independently, avoiding the exponential growth in the number of knots. This ensures computational feasibility as the number of dimensions increases. In the one-dimensional case, there are $K+1$ knots defining the intervals between consecutive segments of the spline. For the remainder of NSF section, we will denote the $i$-th component of $\bm{\theta}$ and $\mathbf{z}$ as $\theta$ and $z$, respectively. This spline is composed of segments that lie between $K+1$ coordinate pairs known as knots $(\theta_{k},z_{k})$, for $k=0,\ldots,K$, which refer to as knots. These knots are organized in a sequence that increases monotonically, starting from $(\theta_{0},z_{0})=(-B,-B)$ and ending at $(\theta_{K},z_{K})=(B,B)$. To define the slope at the internal knots of the spline, $K-1$ arbitrary positive values are assigned, denoted as $\tau^{(k)}$, where $k$ ranges from 1 to $K-1$. The derivatives at the boundaries are set to 1,$\tau^{(0)}$ and $\tau^{(K)}$ are set to one, ensuring the slopes at the endpoints are consistent with the linear tails of the spline.
@@
Let $\phi_{k}$ be the slope of the segment between two consecutive knots: $\phi_{k}=\frac{{z}_{k+1}-{z}_{k}}{{\theta}_{k+1}-{\theta}_{k}}$. Additionally, define $\xi({\theta})$ as the normalized position of ${\theta}$ within the $k$-th interval: $\xi({\theta})=\frac{{\theta}-{\theta}_{k}}{{\theta}_{k+1}-{\theta}_{k}}$. In NSF, the parameters of the rational-quadratic spline (such as the widths and heights of the bins, as well as the derivatives at the knots) are predicted by a neural network. The neural network outputs the parameters that control the spline transformation. Specifically, the networks outputs the widths, heights, and slopes at the internal knots, ensuring that the transformation is invertible and smooth across the bins.
@@
Using these definitions, the transformation $\mathbf{h}({\theta})$ in the $k$-th bin of the spline is given by the ratio of two functions, $\omega_{1}^{(k)}(\xi)$ and $\omega_{2}^{(k)}(\xi)$, expressed as:
@@
where $\omega_{1}^{(k)}(\xi)$ and $\omega_{2}^{(k)}(\xi)$ denote the quadratic polynomials governing the transformation within the $k$-th bin. The parameters of these polynomials, including the slopes and bin heights and width, are determined by the neural network. The rational-quadratic spline is constructed within specified bounds $(-B,B)\times(-B,B)$, with the identity function applied outside this range to ensure numerical stability. The value of $B$ is typically chosen before training to cover the expected range of the input ${\theta}$, ensuring that the transformation remains well-behaved across the entire domain
Given the element-wise and monotonic properties of the rational-quadratic transformation, it is feasible to efficiently compute the logarithm of the absolute value of the determinant of the Jacobian matrix of the transformation. This computation involves summing the logarithms of the derivatives of the function with respect to ${\theta}$. The inverse of the rational-quadratic function, such as $\xi({\theta})$, can be analytically determined by inverting Eq. 29, which is equivalent to finding the roots of a quadratic equation. For a comprehensive derivation, one is referred to [45].
Finally, we put cACLs and cSLs together to build the architecture of our proposed inference network, which is conditional on summary statistics $\hat{\mathbf{u}}$, as shown in Fig. 5. Our design strategy employs a composition of multiple cACLs and cSLs arranged sequentially to approximate a complex posterior. This sequence of layers collectively acts as a deep neural network, with the network parameters denoted by $\bm{\alpha}$. The network structure is characterized by an alternation between cACLs and cSLs, ensuring the necessary depth and complexity for estimating complex posteriors. At each stage of the inference process, the output of a given cACL is channeled as the input to the subsequent cSL, thus creating a chain of transformations. In summary, the inference process is bidirectional: the normalizing direction takes the parameters $\bm{\theta}$ and the observations $\hat{\mathbf{u}}$ to map them to a latent Gaussian variable $\mathbf{z}$, such as $\mathbf{z}=\mathbf{h}_{\bm{\alpha}}(\bm{\theta};\hat{\mathbf{u}})$. Conversely, the generative direction recovers the parameters of interest from the latent space, through the inverse function $\bm{\theta}=\mathbf{h}_{\bm{\alpha}}^{-1}(\mathbf{z};\hat{\mathbf{u}})$.
@@ 3.4 Amortized inference
Unlike traditional Bayesian inference methods, which necessitate repeating the entire inference procedures for each new data set, our proposed likelihood-free approach amortizes the workflow by separating it into a computationally demanding training phase followed by a more cost-efficient inference phase. Fig. 6 depicts the schematic of our proposed method. First, we generate synthetic training data using a forward model and a prior distribution. During training, the synthetic data $\{\bm{\theta}^{(i)},\mathbf{u}^{(i)}_{\text{syn},1:k}\}$, is first processed in batches by the summary network $\mathcal{\bm{F}}_{\bm{\beta}}$, producing the set of summary statistics $\{\hat{\mathbf{u}}^{(i)}\}$ which are then fed into the inference network. The networks are jointly optimized by minimizing the loss function in Eq. 22 with respect to all network parameters. During the inference phase, observed data $\mathbf{u}_{o,1:k}$, is input into the trained networks. The posterior distribution is approximated by drawing samples from the latent distribution and leveraging the inference network.
Figure 6: The workflow of proposed amortized likelihood-free inference

### v2

@@ 3.2 Summary network
The summary network serves as a critical preprocessing step, which converts the $u$ measurements into a fixed, low-dimensional input (summary features) to the conditional NF network. This step is important because likelihood-free methods rely on the summary features capturing sufficient information about the measurements to ensure the inference quality. While summary statistics have traditionally been manually crafted based on domain knowledge, such choices may be unreliable, potentially losing important information and adversely affecting the inference process.
The adoption of a summary network offers several advantages, particularly in reducing the data dimensionality and improving the efficiency of NF training. First, it alleviates the computational burden by compressing high-dimensional raw data into a more manageable form, which enhances both training speed and precision. Secondly, preprocessing through the summary network helps remove redundancies and noise inherent in the raw data, which ensures that only the most relevant information is passed on to the inference stage. Finally, the summary network automates the extraction of informative features from the data, eliminating the need for manual feature selection [24].
The summary network consists of the parameterized transformation
@@
where $\hat{\mathbf{u}}$ represents the learned summary features from the raw data $\mathbf{u}_{o,1:k}$, which will be utilized as input to the conditional NF. The architecture of the summary network is chosen based on the characteristics of the observation data. For instance, bidirectional long short-term memory networks are especially suited for time-series data, enabling the effective handling of sequential measurements that exhibit long memory and nonlinear characteristics. Here we use a one-dimensional convolutional neural network to learn summary features.
@@
$\mathcal{L}(\bm{\alpha},\bm{\beta})=\frac{1}{M}\sum_{i=1}^{M}\left\{\frac{1}{2}\left[\mathbf{h}_{\bm{\alpha}}\left(\bm{\theta}^{(i)};\mathcal{\bm{F}}_{\bm{\beta}}\left(\mathbf{u}_{o,1:k}^{(i)}\right)\right)\right]^{2}-\log\left|\det\left(\frac{\partial\mathbf{h}_{\bm{\alpha}}\left(\bm{\theta}^{(i)};\mathcal{\bm{F}}_{\bm{\beta}}\left(\mathbf{u}^{(i)}_{o,1:k}\right)\right)}{\partial\bm{\theta}}\right)\right|\right\}.$ (23)
@@ 3.3 Inference network
The core element of NFs is invertible transformations. In our study, we employ two families of parameterized invertible transformations: conditional ACLs (cACLs), and conditional SLs (cSLs). In this section, we present an inference network combining cACLs and cSLs that is capable of bidirectional mapping between posterior and latent distributions. This dual mapping capability is critical for enabling the forward transformation as well as the reverse propagation.
3.3.1 Conditional affine coupling layers
Dinh et al. [44] introduced ACLs as a technique to create invertible maps with cheap-to-evaluate Jacobian determinants. Each ACL operates by partitioning the input vector $\bm{\theta}$ into two segments, $\bm{\theta}_{1}$ and $\bm{\theta}_{2}$, and similarly dividing the output vector $\mathbf{z}$ into $\mathbf{z}_{1}$ and $\mathbf{z}_{2}$. This division allows us to define invertible transformations in terms of embedded subnetworks, denoted as $\mathbf{s}_{1}(\cdot)$, $\mathbf{s}_{2}(\cdot)$, $\mathbf{t}_{1}(\cdot)$, and $\mathbf{t}_{2}(\cdot)$. These subnetworks, which can be chosen as arbitrary neural network architectures—like fully connected networks—are not required to be invertible. With the observed summary data $\hat{\mathbf{u}}$ serving as an additional conditioning input, the forward transformations as shown in Fig. 3(a) are formulated as
@@ 3.3.1 Conditional affine coupling layers
Figure 3: Illustration of transformations in cACLs.
The forward transformation can be readily inverted. Given $\mathbf{z}=(\mathbf{z}_{1},\mathbf{z}_{2})$, the inversion is accomplished by inversely transforming through cACLs. The inverse transformation, as shown in Fig. 3(b), is formulated as follows
@@
$\bm{\theta}_{1}=(\mathbf{z}_{1}-\mathbf{t}_{2}(\mathbf{\bm{\theta}}_{2},\hat{\mathbf{u}}))\odot\exp(-\mathbf{s}_{2}(\mathbf{z}_{2},\hat{\mathbf{u}})).$ (27)
Eqs. 2427 are essential in guaranteeing the tractability of the Jacobian determinant of the transformation, like in Eq. 14. Within an ACL, the Jacobian matrix is usually triangular, leading to cost-effective evaluation of its determinant.
@@
$\frac{\partial(\bm{\theta}_{1},\bm{\theta}_{2})}{\partial(\mathbf{z}_{1},\mathbf{z}_{2})}=\begin{pmatrix}\frac{\partial\bm{\theta}_{1}}{\partial\mathbf{z}_{1}}&\frac{\partial\bm{\theta}_{1}}{\partial\mathbf{z}_{2}}\\ \frac{\partial\bm{\theta}_{2}}{\partial\mathbf{z}_{1}}&\frac{\partial\bm{\theta}_{2}}{\partial\mathbf{z}_{2}}\end{pmatrix}=\begin{pmatrix}\mathbf{I}&0\\ \frac{\partial\bm{\theta}_{2}}{\partial\mathbf{z}_{1}}&\frac{\partial\bm{\theta}_{2}}{\partial\mathbf{z}_{2}}.\end{pmatrix}$ (28)
@@
In this block triangular structure, the upper left block is the identity matrix $\mathbf{I}$ because $\bm{\theta}_{1}$ is transformed independently of $\mathbf{z}_{2}$. The 0 in the top-right block reflects the fact that $\bm{\theta}_{1}$ is independent of $\mathbf{z}_{2}$. The lower left block captures the dependence of $\bm{\theta}_{2}$ on $\mathbf{z}_{1}$, while the lower right block represents the transformation of $\bm{\theta}_{2}$ with respect to $\mathbf{z}_{2}$. The determinant of a block triangular matrix is simply the product of the determinants of the diagonal blocks. Because the upper left block is the identity matrix, its determinant is 1, and the overall determinant of the Jacobian matrix is determined by the diagonal terms.
3.3.2 Conditional spline layers
@@
Rational-quadratic neural spline flows are designed to model complex distributions with high efficiency and flexibility [45]. These flow models employ monotonic rational-quadratic SLs as the fundamental building block for developing invertible functions. Each bin within these layers is characterized by a monotonically increasing rational-quadratic function. Rational-quadratic functions, defined as the ratio of two quadratic polynomials, are not only differentiable but also analytically invertible when restricted to monotonic segments. The knot points refer to the boundaries between consecutive bins in the spline. At each knot point, the spline must satisfy certain conditions, such as continuity in both the function’s value and its derivative, ensuring smooth transitions between the segments. SLs provides more flexibility than traditional quadratic functions because they allow direct control over the values and derivatives at each knot point. A monotonic rational-quadratic spline is designed to map an interval of $[-B,B]$ onto itself. Outside of this range, the transformation defaults to an identity function, thereby introducing linear “tails” to the curve.
Figure 4: Monotonic rational-quadratic transformation (spline layer).
@@
Fig. 4 illustrates a monotonic rational-quadratic transformation, also known as a spline layer, which is conditioned on the observed data $\hat{\mathbf{u}}$. Splines are applied component-wise, meaning that each dimension is transformed independently, avoiding the exponential growth in the number of knots. This ensures computational feasibility as the number of dimensions increases. In the one-dimensional case, there are $K+1$ knots defining the intervals between consecutive segments of the spline. For the remainder of this section, we will denote the $i$th component of $\bm{\theta}$ and $\mathbf{z}$ as $\theta$ and $z$, respectively. This spline is composed of segments that lie between $K+1$ coordinate pairs known as knots $(\theta_{k},z_{k})$, for $k=0,\ldots,K$. These knots are organized in a sequence that increases monotonically, starting from $(\theta_{0},z_{0})=(-B,-B)$ and ending at $(\theta_{K},z_{K})=(B,B)$. To define the slope at the internal knots of the spline, $K-1$ arbitrary positive values are assigned and denoted as $\tau^{(k)}$, where $k$ ranges from 1 to $K-1$. The derivatives at the boundaries are set to 1,$\tau^{(0)}$ and $\tau^{(K)}$ are set to one, ensuring the slopes at the endpoints are consistent with the linear tails of the spline.
@@
Let $\phi_{k}$ be the slope of the segment between two consecutive knots: $\phi_{k}=\frac{{z}_{k+1}-{z}_{k}}{{\theta}_{k+1}-{\theta}_{k}}$. Additionally, define $\xi({\theta})$ as the normalized position of ${\theta}$ within the $k$th interval: $\xi({\theta})=\frac{{\theta}-{\theta}_{k}}{{\theta}_{k+1}-{\theta}_{k}}$. In SLs, the parameters of the rational-quadratic spline (such as the widths and heights of the bins, as well as the derivatives at the knots) are predicted by a neural network. The neural network outputs the parameters that control the spline transformation. Specifically, the networks outputs the widths, heights, and slopes at the internal knots, ensuring that the transformation is invertible and smooth across the bins.
@@
Using these definitions, the transformation $\mathbf{h}({\theta})$ in the $k$th bin of the spline is given by the ratio of two functions, $\omega_{1}^{(k)}(\xi)$ and $\omega_{2}^{(k)}(\xi)$, expressed as:
@@
where $\omega_{1}^{(k)}(\xi)$ and $\omega_{2}^{(k)}(\xi)$ denote the quadratic polynomials governing the transformation within the $k$th bin. The parameters of these polynomials, including the slopes and bin heights and width, are determined by the neural network. The rational-quadratic spline is constructed within specified bounds $(-B,B)\times(-B,B)$, with the identity function applied outside this range to ensure numerical stability. The value of $B$ is typically chosen before training to cover the expected range of the input ${\theta}$, ensuring that the transformation remains well-behaved across the entire domain
Given the element-wise and monotonic properties of the rational-quadratic transformation, it is feasible to efficiently compute the logarithm of the absolute value of the determinant of the Jacobian matrix of the transformation. This computation involves summing the logarithms of the derivatives of the function with respect to ${\theta}$. The inverse of the rational-quadratic function, such as $\xi({\theta})$, can be analytically determined by inverting Eq. 29, which is equivalent to finding the roots of a quadratic equation [45].
Finally, we put cACLs and cSLs together to build the architecture of our proposed inference network, which is conditional on summary features $\hat{\mathbf{u}}$, as shown in Fig. 5. Our design strategy employs a composition of multiple cACLs and cSLs arranged sequentially to approximate a complex posterior. This sequence of layers collectively acts as a deep neural network, with the network parameters denoted by $\bm{\alpha}$. The network structure is characterized by an alternation between cACLs and cSLs, ensuring the necessary depth and complexity for estimating complex posteriors. At each stage of the inference process, the output of a given cACL is channeled as the input to the subsequent cSL, thus creating a chain of transformations. In summary, the inference process is bidirectional: the normalizing direction takes the parameters $\bm{\theta}$ and the observations $\hat{\mathbf{u}}$ to map them to a latent Gaussian variable $\mathbf{z}$, such as $\mathbf{z}=\mathbf{h}_{\bm{\alpha}}(\bm{\theta};\hat{\mathbf{u}})$. Conversely, the generative direction recovers the parameters of interest from the latent space through the inverse function $\bm{\theta}=\mathbf{h}_{\bm{\alpha}}^{-1}(\mathbf{z};\hat{\mathbf{u}})$.
@@ 3.4 Amortized inference
Unlike traditional Bayesian inference methods, which necessitate repeating the entire inference procedures for each new dataset, our proposed likelihood-free approach amortizes the workflow by separating it into a computationally demanding training phase followed by a more cost-efficient inference phase. Fig. 6 depicts the schematic of our proposed method. First, we generate synthetic training data using a forward model and a prior distribution. During training, the synthetic data $\{\bm{\theta}^{(i)},\mathbf{u}^{(i)}_{\text{syn},1:k}\}$, is first processed in batches by the summary network $\mathcal{\bm{F}}_{\bm{\beta}}$, producing the set of summary features $\{\hat{\mathbf{u}}^{(i)}\}$ that are then fed into the inference network. The networks are jointly optimized by minimizing the loss function in Eq. 22 with respect to all network parameters. During the inference phase, observed data $\mathbf{u}_{o,1:k}$, is input into the trained networks. The posterior distribution is approximated by drawing samples from the latent distribution and leveraging the inference network.
Figure 6: The workflow of proposed amortized likelihood-free inference.